You are here

ENR CTIO Blanco 4-m 2014-11-29

Primary tabs

Date: 
Saturday, November 29, 2014
Telescope: 
CTIO Blanco 4-m
Submitter: 
Chris Smith
Submitter Email: 
Observer Support: 
Rodrigo Hernandez
Night Assistant: 
Claudio Aguilera
Also Present: 
Javier Rojas, Humberto Orrego, Lucy Clerkin, Nicokay Kuropatkin
Problems: 

1) Clouds
2) Strange spike in pressure in 1r line, accompanied by frost on line for short time; watched carefully, but no additional problems
3) primary mirror significantly warmer than ambient

Comments: 
*Observing Team*
 
_DES Observers:_ Nikolay Kuropatkin (Obs1), Chris Smith (Obs2), Lucy Clerkin (RM) 
_Telops:_ Rodrigo Hernandez (Obs Support), Claudio Aguilera (Night Support)
_Remote Support:_ ---
 
*Observing Plan*
 
# Standard bias, standard flats
# evening standard stars 
# run ObsTac queue
# end of the night standard star
# end of night calibrations 
 
*Conditions*
 
# weather in afternoon 17:30 (local): thick cloud cover, 7.2C , wind 7 mph, humidity 65%
# at sunset 20:30 local: clouds over most of sky, clear band to west, 7.1C, wind 6.9 mph, humidity 48%
# at 22:50, clouds have moved enough to allow opening, sky 3/4 clear (cloud band now covering only 1/4 to East)
# 22:50 local: 6.2C, wind 4 mph, humidity 52%
# at start of OCS, 23:10 local: 3/4 sky clear, 6.3C, wind 4 mph, humidity 48%
# NOTE: mirror temp significantly higher than outside temp (about +3C) = bad seeing
# 01:30 local: 7.1C, wind 5 mph, humidity 25%
# 03:10 local: 6.7C, wind 10mph, humidity 38%
# raised windscreen at 03:15, wind speed consistently >9mph
# 05:10 local: 7.2C, wind 7 mph, humidity 30%
 
# weather summary: mostly clear, 6-7C, humidity 30-60%
# Seeing conditions: ~1.5-3.0, highly variable during night
# Wind conditions: wind 5-10 mph
 
*Accomplishments*
 
# Afternoon:  SISPI restarted, zeros done, (YziVRrgu) dome flats done, flats checked OK
# Dome NOT opened before sunset due to threatening clouds
# 22:50 local time: clouds moving out to east, open dome!
# seeing in zero points = 3"... not unexpected as clouds move out
# 23:10 local: start OCS queue; SUPERNOVAE (perhaps clean up observations from rejected frames from last night?)
# SN field E1 (griz), seeing g=2.75", r=2.61", i=2.30", z=2.55" (seeingall)
# SN field E2 (griz), seeing g=2.49", r=2.40", i=2.35", 
# SN field X2 (griz), seeing g=2.69"
# SN field C2 (griz), seeing g=2.60"
# SN field C1 (griz), seeing g=2.45"
# 00:57 local: SWITCH to wide-field survey
# survey exposures: all Y-band (plus three z) through 02:15, while moon up
# survey exposures: g band starting at 02:16 local
# moon set at 02:28 local
# survey exposures almost all g band for rest of night (except one Y and r)
# survey exposures: started Y band at 05:12 local
# End of night: Standards (05:25 local)
## C26202/HST
## SDSS J1048 - 0000
## SDSS J0843 - 0000  
 
 
*Problems tagged as Synopsis_Problem*
 
# Strange spike in 1r nitrogen line pressure, accompanied by slight frost on the 1r metal lines; system seemed to recover without intervention.
 
# Rasicam: off
# DIMM: on
 
 
*Minor Problems and Relevant Information*
 
# primary mirror temp significantly (~3C) higher than outside temp, leading to bad seeing
 
*Notes for Data Management*
 
# clear but bad seeing. 
 
Conditions: 
General Conditions
Comment: 
None.
1st Quarter
Clouds: 
Partially cloudy
Wind: 
Low
Seeing Value: 
0.00
Seeing Variability: 
0
Seeing Source: 
CTIO Seeing monitor
Sky : 
6
2nd Quarter
Clouds: 
Thin Clouds
Wind: 
Low
Seeing Value: 
2.50
Seeing Variability: 
0.4
Seeing Source: 
The Telescope I used
Sky: 
2
3rd Quarter
Clouds: 
Clear
Seeing Value: 
2.70
Seeing Variability: 
0.4
Seeing Source: 
The Telescope I used
Sky: 
1
4th Quarter
Clouds: 
Clear
Seeing Value: 
2.70
Seeing Variability: 
0.4
Seeing Source: 
The Telescope I used
Sky: 
0
Program: 
Proposal: 
2012B-0001
Target of Oportunity: 
No
Time Observed: 
5.5
Time Engineering: 
0.0
Time lost technical: 
0.0
Time lost weather: 
2.5
Time lost other: 
0.0
Description: 
Dark Energy Survey
Primary Investigator: 
Frieman
Institution: 
DES
Instrument: 
CTIO 4m DECam
Organization: 
NOAO
Total Program Time: 
8
Total Times Time Observed Time Engineering Time lost technical Time lost weather Time lost other Total Program Time
2014-11-29 5.5 0.0 0.0 2.5 0.0 8